OpenStudy (anonymous):

Which of the following is the same as 10 4?  4 x 4 x 4 x 4 x 4 x 4 x 4 x 4 x 4 x 4  40 x 40 x 40 x 40  10 x 4  10 x 10 x 10 x 10

OpenStudy (dqrodgers1998):

10 x 4 and 4 x 4 x 4 x 4 x 4 x 4 x 4 x 4

OpenStudy (dqrodgers1998):

both equal 40

OpenStudy (anonymous):

10x10x10x10

OpenStudy (anonymous):

are you talking about 10^4?

OpenStudy (anonymous):

d

OpenStudy (dqrodgers1998):

oh i did it wrong d is right

OpenStudy (arabpride):

10 x 10 x 10 x 10 = 10^4

OpenStudy (anonymous):

D

OpenStudy (anonymous):

its basically saying 10 for times so if it said 4 to 10 that would make it A but in this case its asking 10 4 times so its D

OpenStudy (skullpatrol):

$$\Huge b^n=\underbrace{b*b*b*...*b}$$ $$\Huge\text{n -factors}$$

OpenStudy (anonymous):

oh my bad sorry wrong solution

OpenStudy (anonymous):

its "D"

OpenStudy (skullpatrol):

$$\Huge 10^4=10*10*10*10$$

OpenStudy (skullpatrol):